OpenStudy (anonymous):

Becky is 18 and would like to buy a house when she is 36. What is the discount factor for today’s prices if the housing values increase 6% per year?

ganeshie8 (ganeshie8):

\((1+.06)^{36-18}\) ?

OpenStudy (anonymous):

that wouldnt be one of my answer choices...

OpenStudy (anonymous):

i tried it

ganeshie8 (ganeshie8):

hmm i think thats the answer

ganeshie8 (ganeshie8):

scratch that

ganeshie8 (ganeshie8):

here question is actually asking to find the factor by which you multiply the future value, to get the current value, right ?

OpenStudy (anonymous):

Well, i know the answer has to be a %

OpenStudy (anonymous):

so i suppose thats a yes

ganeshie8 (ganeshie8):

ok.. then we should do a inverse : 1/(1+.06)^18 = 0.350343791

OpenStudy (anonymous):

i just did that and mine came out to 2.84 :/

ganeshie8 (ganeshie8):

inverse it..

OpenStudy (anonymous):

but if i multiply your answer by 100, i get one of my answer choices which is 35%

ganeshie8 (ganeshie8):

yeah (future value) = (current value) * (1+.06)^18 right ?

OpenStudy (anonymous):

No, i only multiplied by 100 to get a % instead of a decimal

ganeshie8 (ganeshie8):

(current value) = (1/(1+.06)^18) * (future value)

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

this our discount factor then (1/(1+.06)^18)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw ^_^

OpenStudy (anonymous):

do you maybe know how to calculate this Rudy has been awarded some money in a settlement. He has the option to take a lump sum payment of $200,000 or get paid an annuity of $1,000 per month for the next 25 years. Which is the better deal for Rudy, and by how much, assuming the growth rate of the economy is 2.75% per year?

ganeshie8 (ganeshie8):

noo im scared of annuity... lol they are tough problems :\ post it freshly... someone wil help u gud luck :)

OpenStudy (anonymous):

i have to use FV=PV(1+i)^nt backwards but idk how. But thank you

ganeshie8 (ganeshie8):

i think you have to use the annuity formula... i know how to do this. but i dont know why it works :( i only know only the formula...

OpenStudy (anonymous):

oh wait, i have to divide the result of (1+i) on both sides

ganeshie8 (ganeshie8):

here, i = r/n ?

OpenStudy (anonymous):

35.0%